OpenStudy (anonymous):

How would I factor x^4 - 13x^2 + 36 = 0 ? I have NO idea how to do this.

OpenStudy (anonymous):

would I get (x^2 -4x)(x^2 -9x)=0 ?

OpenStudy (anonymous):

x^4 - 13x^2 + 36 = 0 x^4 - 9x^2-4x^2 + 36 = 0 x^2(x^2-9)-4(x^2-9)=0 (x^2 -4x)(x^2 -9x)=0